21 January 2008

RHIC Running:  Gruesome weekend.  The accelerator is not behaving in a reproducible manner.  There are many problems, including instabilities that cause the emittance of the Yellow beam to blow up.  Emittance at injection to the Booster is about 5p mm-mr, but by the time the beam reaches RHIC it is about 10p mm-mr and growing.  Under some circumstances the emittance growth is much greater, only we are not clear about what those circumstances are.  Power supply problems in RHIC (bo11-qd3-ps) required an access to replace the faulty supply.  Controls had to restart two RHIC managers. 

Angelika will do a vernier scan at the end of the morning store, around 11:00AM.

The AGS L-10 cavity tripped off again, as it has been giving problems all weekend.  When the L-10 drops out, the bunch intensity is low, and sometimes "baby bunches" develop in deuterons. 

In addition, Booster-to-AGS synchro is having problems during Au set-up.

Development of polarized protons has progressed well.  Protons have been brought out of the AGS.  No polarization measurement yet in the AGS, but that should be accomplished soon.


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020